# Multi-Source Urban Traffic Flow Forecasting with Drone and Loop Detector Data

### v1

@@ Abstract
Traffic forecasting is a fundamental task in transportation research, however the scope of current research has mainly focused on a single data modality of loop detectors. Recently, the advances in Artificial Intelligence and drone technologies have made possible novel solutions for efficient, accurate and flexible aerial observations of urban traffic. As a promising traffic monitoring approach, drone-captured data can create an accurate multi-sensor mobility observatory for large-scale urban networks, when combined with existing infrastructure. Therefore, this paper investigates the problem of multi-source traffic speed prediction, simultaneously using drone and loop detector data. A simple yet effective graph-based model HiMSNet is proposed to integrate multiple data modalities and learn spatio-temporal correlations. Detailed analysis shows that predicting accurate segment-level speed is more challenging than the regional speed, especially under high-demand scenarios with heavier congestions and varying traffic dynamics. Utilizing both drone and loop detector data, the prediction accuracy can be improved compared to single-modality cases, when the sensors have lower coverages and are subject to noise. Our simulation study based on vehicle trajectories in a real urban road network has highlighted the added value of integrating drones in traffic forecasting and monitoring.
@@ I Introduction
Owing to the limitations of data availability, many deep learning methods are built upon highway loop detector data, e.g., METR-LA and PEMS-Bay [8]. However, urban traffic forecasting presents additional challenges due to the dense and complex road network structure. Recently, Unmanned Aerial Vehicles (UAVs) or drones are emerging as a novel solution to information collection in ITS [9], as they can flexibly fly over an area of interest and record high-quality videos. At the same time, the growing power of Computer Vision, Artificial Intelligence and computational resources have made it possible to automatically and efficiently extract object information from the captured videos. These advantages have empowered drones to gather rich and accurate information from a large spatial region, making them a preferable choice for urban traffic monitoring.
@@
Specifically, the pNEUMA experiments have started the understanding of urban congestion evolution through the analysis of video data captured by a swarm of drones over a large area in the city center of Athens [10]. The drone-captured videos present clear views of the monitored areas, and most vehicles can be identified by object detection and tracking methods [11]. Detailed trajectories collected from drones can significantly improve the understanding of various traffic phenomena, including emissions [12], multi-modal MFDs [13], path flow estimation [14], car following [15], traffic signal safety [16] and parking monitoring [17]. With vehicle trajectories, traffic variables can be accurately computed for all visible roads in the field of view (FOV) [18], resulting in a great advantage compared to grounded sensors, e.g., loop detectors and CCTV cameras, whose scopes are limited to the roads with installed facilities.
@@
The high-quality data from drones can provide solid grounds for traffic state prediction, which will be an additional asset to the existing traffic forecasting methods based on loop detector data. However, few studies have attempted to address the traffic forecasting problem using joint observatories from drones and other information sources. Therefore, this paper aims to investigate a promising paradigm of urban traffic forecasting, where drone-measured traffic speed data are combined with loop detector data to provide high-quality speed predictions.
@@
In summary, the main contributions are as follows. First, this paper introduces a flexible model architecture HiMSNet for road segment-level speed prediction. It consists of a local module for data fusion and a global module for spatial dependencies learning, and it is further extended with an additional branch for regional speed prediction. Second, A well-calibrated multi-source dataset, i.e., SimBarca, is created, offering traffic speed data from drones and loop detectors in the city center of Barcelona. Using the vehicle trajectories simulated in AIMSUN microsimulation software, SimBarca gathers statistics for the prediction tasks of segment-level and regional speeds. Third, Extensive experiments and analysis are provided to evaluate HiMSNet on SimBarca, providing an example for evaluating multi-source urban traffic forecasting. Finally, The methods, data and evaluations in this work are openly shared with the community to facilitate future research. (Link will be provided upon acceptance)
@@ II-A GNN-Based Methods
Since the road network has a natural graph structure, Graph Neural Networks have been extensively explored in traffic forecasting [21]. In the modern era of deep learning, the Diffusion Convolutional Recurrent Neural Network (DCRNN) [8] is one of the pioneering works to utilize a graph-based model in this area. In DCRNN, the temporal dependencies are learned by a Recurrent Neural Network (RNN), whose layer has an embedded diffusion convolution module to model the spatial dependencies by a diffusion process in the sensor graph. A network block in Spatio-temporal Graph Convolutional Neural Networks (STGCN) [22] applies a convolutional operation over the time dimension, and then propagates the information over the spatial graph structure with Graph Convolution. Following a similar design philosophy, the Temporal Graph Convolutional Network (T-GCN) [23] propagates neighborhood information over the graph at each time step and then applies an RNN to learn the temporal dependencies.
@@ II-B Transformer-Based Methods
Yan et al. [34] use a transformer model to learn the spatial dependencies, and employ an LSTM [35] for temporal learning. The adjacency matrix is used in transformer decoder part to enforce the spatial constraints. Xu et al. [36] propose to learn temporal relations with transformers, and fuse the output of transformer and Graph Neural Network (GNN) to obtain a joint representation of the spatial relations. In this way, the model can make use of the flexibility of pair-wise correlations in the transformer model, and also follow the spatial dependencies of the graph structure. Similar to GMAN [29], the Meta Graph Transformer (MGT) [37] designs a cross-attention between encoder and decoder, which allows the model to capture the relations between historical and future time steps. To enhance the graph structure, MGT also encoded multiple weight matrices based on geographical distance, historical traffic flow and vehicle counts. Shao et al. [38] adapted the method from Masked Auto-Encoder [39], and propose to obtain a better data representation by a pretraining phase on multiple datasets, where a transformer model is trained to complete the missing values with only partial inputs. PDFormer [40] clusters segments of traffic time series into sets of representative patterns, and utilize them to identify similar patterns in new input sequences when making predictions.
Since traffic forecasting datasets usually do not have a big scale, building a complicated model is not always necessary. For example, STID [41] achieves similar accuracy as PDFormer by using MLPs only. A later study on general time series forecasting has also shown that linear models can achieve similar performance as transformer models [42] if properly tuned. Compared to existing literature, this study focuses on an urban area with denser road networks and highly dynamic traffic. At the same time, multiple data sources are utilized together to provide more accurate and robust predictions. In previous studies, the research scope is often the traffic flow prediction on highways using data from loop detectors only. This work aims to provide a simple yet effective model along with comprehensive evaluations to lay the foundation for future research. Therefore, we refrain from using a complex model design and keep the implementation concise.
@@ III Methodology
This section will first introduce the formulation of multi-source traffic forecasting. Then a baseline model for multi-source traffic forecasting will be presented, and the final part will describe the dataset creation method using vehicle trajectories.
@@ III-A Multi-Source Traffic Forecasting
where the time steps $t_{i}$ are chronologically sorted. Such a time series has its own evolution patterns over time, and at the same time is correlated with time series from other locations, since they belong to the same transportation network $\mathcal{G}$. Besides, since different observation methods reveal different aspects of the traffic states, i.e., different sensor coverages, the inter-modality correlation is also important in multi-source traffic forecasting. As a result, if there are $M$ observation methods and $P$ locations in total, a Multi-Source Time Series (MSTS) is the collection of time series from all modalities and locations, which comprehensively describe the available traffic information in the transportation network.
$\mathcal{S}=\left\{\mathcal{S}^{i,j}\right\},\text{where}~{}i\in[1,2,...,M],~{}j\in[1,2,...,P].$ (3)
@@
where $l$ is a distance function to measure the differences between predictions and labels. The exact form of the label $\mathbf{y}_{i}$ and loss $l$ will depend on the task (e.g., forecasting, interpolation and imputation), and more details will be provided in Section IV-B, where the experiment details will be elaborated. Besides, the notations of time series sequences and dataset are inspired by [43].
@@ III-B Model Architecture
Figure 1 shows the structure of our baseline model Hierarchical Multi-Source Network (HiMSNet) for traffic forecasting. This architecture consists of a Global Message Exchange (GME) module to learn the spatial correlations between different locations, and a Local State Evolution (LSE) module to handle temporal dependencies and multiple data modalities at each location. It is a simple and straightforward baseline model that addresses the spatial, temporal and modality correlations in multi-source traffic forecasting.
Following the road network, the input data can be naturally organized as a graph structure, where a node corresponds to a road segment, and an edge indicates the connection between two road segments. In the LSE module, the temporal encoder $\mathcal{H}_{e}$ will take $\mathcal{S}^{*,p}$, i.e., the time series from all observation methods at location $p$ (modalities marked in colors), and separately encode the time series of each modality into a unified representation $\mathbf{z}_{p}$. Then, $\mathbf{z}_{p}$ will go through a message encoding layer $\mathcal{R}_{1}$ and become a message to be shared with other locations. The GME module will then take the messages from all locations ($\mathbf{Z}$), use its encoder $\mathcal{F}_{e}$ and decoder $\mathcal{F}_{d}$ to compute the new feature with exchanged information ($\mathbf{Q}$). The message decoding layer $\mathcal{R}_{2}$ will take the element in $\mathbf{Q}$ corresponding to location $p$, and concatenate it together with the elements in $\mathbf{z}_{p}$ to form a joint feature $\mathbf{q}_{p}$, which contains the information from the location $p$ itself and its vicinity. Finally, the temporal decoder $\mathcal{H}_{d}$ will take $\mathbf{q}_{p}$ and predict the future traffic states $\hat{\mathbf{y}}_{p}$.
@@
Figure 1: The model structure of HiMSNet. The model consists of a global layer and a local layer. The local layer processes the time series from the sensors that are monitoring a road segment, and obtains a joint feature for each road segment. The global layer will initiate message exchanges across the road network, and enhance the local joint features with vicinity information.
Figure 2: Baseline branch for regional prediction. This branch gathers the features of all road segments in a region and predicts the regional traffic.
@@
HiMSNet is flexible and can be easily extended with an additional branch (Figure 2) for regional traffic forecasting, which can be helpful in regional traffic control. Given a defined region $I$, the regional branch will take the average of joint feature $\mathbf{q}_{p}$ for all $p\in I$ as the regional representation $\mathbf{q}_{I}$, and predict the regional average traffic state $\hat{\mathbf{y}}_{I}$ with the prediction module $\mathcal{H}_{I}$.

### v2

@@ Abstract
Traffic forecasting is a fundamental task in transportation research, however the scope of current research has mainly focused on a single data modality of loop detectors. Recently, the advances in Artificial Intelligence and drone technologies have made possible novel solutions to efficient, accurate and flexible aerial observations of urban traffic. As a promising traffic monitoring approach, drones can create an accurate multi-sensor mobility observatory for large-scale urban networks, when combined with existing infrastructure. Therefore, this paper investigates urban traffic prediction from a novel perspective, where multiple input sources, i.e., loop detectors and drones, are utilized to predict both the individual traffic of all road segments and the regional traffic. A simple yet effective graph-based model, Hierarchical Multi-Source Neural Network (HiMSNet), is proposed to integrate multiple data modalities and learn spatio-temporal correlations. Detailed analysis shows that predicting accurate segment-level speed is more challenging than the regional speed, especially under high-demand scenarios with heavier congestions and varying traffic dynamics. Utilizing both drone and loop detector data, the prediction accuracy can be improved compared to single-modality cases, when the sensors have lower coverages and are subject to noise. Our simulation study based on vehicle trajectories in a real urban road network has highlighted the value of integrating drones in traffic forecasting and monitoring.
@@ I Introduction
Owing to the limitations of data availability, many deep learning methods are built upon highway loop detector data, e.g., METR-LA and PEMS-Bay [8]. However, urban traffic forecasting presents additional challenges due to the dense and complex road network structure. Recently, Unmanned Aerial Vehicles (UAVs) or drones are emerging as a novel information-gathering solution in ITS [9], as they can flexibly fly over an area of interest and record high-quality videos. At the same time, the growing power of Computer Vision, Artificial Intelligence and computational resources have made it possible to automatically and efficiently extract vehicle locations from the captured videos. These advantages have empowered drones to collect rich and accurate information from a large spatial region, making them a preferable choice for urban traffic monitoring.
@@
Specifically, Barmpounakis et al. [10] have conducted a large-scale experiment in Athens, where a swarm of drones was deployed to monitor traffic in the city center. The drone-captured videos can support in-depth analysis and understanding of urban traffic, because they present clear views of the monitored areas, and most vehicles can be identified by object detection and tracking methods [11]. With the detailed vehicle trajectories, the understanding of various traffic phenomena can be significantly improved, including emissions [12], multi-modal Macroscopic Fundamental Diagrams (MFD) [13], path flow estimation [14], car following [15], traffic signal safety [16] and parking monitoring [17]. With vehicle trajectories, traffic variables can be accurately computed for all visible roads in the field of view (FOV) [18], resulting in a great advantage compared to grounded sensors, e.g., loop detectors and video suveillance cameras, whose scopes are limited to the roads with installed facilities.
@@
The high-quality data from drones can provide solid grounds for traffic state prediction, which will be an additional asset to the existing traffic forecasting methods based on loop detector data. However, few studies have attempted to address the traffic forecasting problem using joint observatories from drones and other information sources. Besides, little attention has been paid to deep learning-based regional traffic predictions, which is an essential task for urban traffic management. Therefore, this paper aims to investigate a promising paradigm of multi-source urban traffic forecasting, where drone-measured traffic are combined with loop detector data to provide high-quality predictions for both individual road segments and urban regions at the same time.
@@
In summary, the main contributions are as follows. First, this paper introduces a flexible deep learning model HiMSNet to jointly handle road segment-level and regional traffic prediction, using traffic measurements of simulated drones and loop detectors. Second, A well-calibrated multi-source dataset SimBarca is created, offering simulated traffic speed data of drones and loop detectors in Barcelona. Using the simulated trajectories of all vehicles in a dense urban road network, SimBarca introduces the prediction tasks of segment-level and regional traffic speeds. Third, extensive experiments and analysis are provided to evaluate HiMSNet on SimBarca, highlighting the necessity of drone data in accurate urban traffic prediction. Finally, the data, model and evaluations in this work are openly shared with the community to facilitate future research. (Link will be provided upon acceptance)
@@ II-A GNN-Based Methods
Since the road network has a natural graph structure, Graph Neural Networks (GNN) have been extensively explored in traffic forecasting [21]. In a GNN-based model, nodes usually represent sensors (e.g., loop detectors) at various locations, and the edges correspond the road connections between various locations. The model will gather information from the graph neighbors of each node, which are often the other sensors within a dinstance. As a result, GNNs can pass messages according to the graph structure and thus learn spatial relations.
In the modern era of deep learning, the Diffusion Convolutional Recurrent Neural Network (DCRNN) [8] is one of the pioneering works to utilize a graph-based model in this area. In DCRNN, the temporal dependencies are learned by a Recurrent Neural Network (RNN), whose layer has an embedded diffusion convolution module to model the spatial dependencies by a diffusion process in the sensor graph. A network block in Spatio-temporal Graph Convolutional Neural Networks (STGCN) [22] applies a convolutional operation over the time dimension, and then propagates the information over the spatial graph structure with Graph Convolution. Following a similar design philosophy, the Temporal Graph Convolutional Network (T-GCN) [23] propagates neighborhood information over the graph at each time step and then applies an RNN to learn the temporal dependencies.
@@ II-B Transformer-Based Methods
Yan et al. [34] use a transformer model to learn the spatial dependencies, and employ an Long Short-Term Memory Network (LSTM) [35] for temporal learning. The adjacency matrix is used in transformer decoder part to enforce the spatial constraints. Xu et al. [36] propose to learn temporal relations with transformers, and fuse the output of transformer and Graph Neural Network (GNN) to obtain a joint representation of the spatial relations. In this way, the model can make use of the flexibility of pair-wise correlations in the transformer model, and also follow the spatial dependencies of the graph structure. Similar to GMAN [29], the Meta Graph Transformer (MGT) [37] designs a cross-attention between encoder and decoder, which allows the model to capture the relations between historical and future time steps. To enhance the graph structure, MGT also encoded multiple weight matrices based on geographical distance, historical traffic flow and vehicle counts. Shao et al. [38] adapted the method from Masked Auto-Encoder [39], and propose to obtain a better data representation by a pretraining phase on multiple datasets, where a transformer model is trained to complete the missing values with only partial inputs. PDFormer [40] clusters segments of traffic time series into sets of representative patterns, and utilize them to identify similar patterns in new input sequences when making predictions.
II-C Research Gap and Contribution
Although traffic forecasting has been widely studied, the scope of existing research has been centered on predicting loop detector-measured highway traffic, due to the limitation of data availability. Aside from single-sensor traffic prediction, some recent research works have explored using multiple data sources to estimate urban traffic states using data combination [41, 42, 43], probabilistic filtering [44, 45], model ensemble [46] and deep feature fusion [47]. However, the scope of these research works is limited to one or a few roads.
In general, traffic forecasting datasets do not have big scales, and thus building complicated models is not always necessary. For example, STID [48] achieves similar accuracy as the transformer-based PDFormer [40] by using Multi-Layer Perceptrons (MLPs) only. A later study on time series forecasting has also shown that linear models can achieve similar performance as transformer models [49] if properly tuned. Therefore, this work will refrain from using a complex model design and will keep the implementation concise, as the primary aim is to provide a simple yet effective model along with comprehensive evaluations to build a solid foundation for future research.
Compared to existing literature, this study expands the research scope in problem scale, data quality and prediction tasks. First, this work aims to predict the traffic of all roads in an urban area with dense connections and highly dynamic traffic. Secondly, the drone-measured traffic can accurately represent the traffic states of complete roads, instead of a short range around the installed loop detectors. Finally, regional traffic prediction, has been made possible provided that the drones can coverage a reasonable percentage of the area.
@@ III Methodology
This section will first introduce the formulation of multi-source traffic forecasting (notation inspired by [50]). Then a baseline model for multi-source traffic forecasting will be presented, and the final part will describe the dataset creation method using vehicle trajectories.
@@ III-A Multi-Source Traffic Forecasting
where the time steps $(t_{0},t_{1},...,t_{n})$ are chronologically sorted. Such a time series has its own evolution patterns over time, and at the same time is correlated with time series from other locations, since they belong to the same transportation network $\mathcal{G}$. Besides, since different observation methods reveal different aspects of the traffic states, i.e., different sensor coverages, the inter-modality correlation is also important in multi-source traffic forecasting. As a result, if there are $M$ observation methods and $P$ locations in total, a Multi-Source Time Series (MSTS) is the collection of time series from all modalities and locations, which comprehensively describe the available traffic information in the transportation network.
$\mathcal{S}=\left\{\mathcal{S}^{m,p}\right\},\text{where}~{}m\in[1,2,...,M],~{}p\in[1,2,...,P].$ (3)
@@
Following the same notation convention, $\mathcal{S}_{k}^{m,p}$ will be the traffic variable measured by method $m$ at location $p$ in the $k$-th sample. And the details for creating the dataset and the samples will be presented in Section III-C.
@@
where $l$ is a distance function to measure the differences between predictions and labels. Specifically, both regional and segment-level speed prediction use the Mean Absolute Error (MAE, as detailed in Equation 11) as the loss function. In case no vehicle is observed, the traffic variable is marked as missing, and will be ignored in loss calculation and performance evaluation. The overall training loss $l$ for a batch is a sum of the segment-level loss $l_{seg}$ and the regional loss $l_{reg}$:
$l=l_{seg}+l_{reg}$ (6)
@@ III-B1 Overall Pipeline
Figure 1 shows the structure of our baseline model Hierarchical Multi-Source Network (HiMSNet) for traffic forecasting. This architecture consists of a Global Message Exchange (GME) module to learn the spatial correlations between different locations, and a Local State Evolution (LSE) module to handle temporal dependencies and multiple data modalities at each location. It is a simple and straightforward model that addresses the spatial, temporal and modality correlations in multi-source traffic forecasting.
@@
Figure 1: The two-layer model structure of HiMSNet. The local layer processes all available time series data of a road segment, and concatenates the feature vectors of different data modalities into a joint feature. The global layer exchanges messages across the road network, and enhance the local joint features with vicinity information.
Following the road network, the input data can be naturally organized as a graph structure, where a node corresponds to a road segment, and an edge indicates the connection between two road segments. Before feeding the data into the deep learning model, all data modalities are normalized to have zero mean and unit variance. In the LSE module, the temporal encoder $\mathcal{H}_{e}$ will take $\mathcal{S}^{*,p}$, i.e., the time series from all observation methods at location $p$ (modalities marked in colors), and separately encode the time series of each modality into a unified representation $\mathbf{z}_{p}$ [^1] An encoder is a neural network module that transforms the input data to a hidden state, and a decoder computes desired outputs from the hidden state. The LSE module will then concatenate the representations of all modalities to form a joint feature $\mathbf{z}_{p}$ for location $p$. Then, $\mathbf{z}_{p}$ will go through a message encoding layer $\mathcal{R}_{1}$ and become a message to be shared with other locations. The GME module will then take the messages from all locations ($\mathbf{Z}$), use its encoder $\mathcal{F}_{e}$ and decoder $\mathcal{F}_{d}$ to compute the new feature with exchanged information ($\mathbf{Q}$). The message decoding layer $\mathcal{R}_{2}$ will take the element in $\mathbf{Q}$ corresponding to location $p$, and concatenate it together with the elements in $\mathbf{z}_{p}$ to form a joint feature $\mathbf{q}_{p}$, which contains the information from the location $p$ itself and its vicinity. Finally, the temporal decoder $\mathcal{H}_{d}$ will take $\mathbf{q}_{p}$ and predict the future traffic states $\hat{\mathbf{y}}_{p}$.
Figure 2: Regional prediction branch. This branch groups the segment-level features by spatial regions (Figure 9), averages the features by group, and use them to predict regional traffic.
@@
HiMSNet is flexible and can be easily extended with an additional branch (Figure 2) for regional traffic forecasting, which can be helpful in regional traffic control. Given a defined region $I$, the regional branch will take the average of joint feature $\mathbf{q}_{p}$ for all $p\in I$ as the regional representation $\mathbf{q}_{I}$, and predict the regional average traffic state $\hat{\mathbf{y}}_{I}$ with the prediction module $\mathcal{H}_{I}$. The components related to this branch are in parallel with the right part of LSE module in Figure 1, i.e., $\mathbf{q}_{p}$, $\mathcal{H}_{d}$ and $\hat{\mathbf{y}}_{p}$, and regional features are averaged from the segment-level features belonging to the region.
III-B2 Building Blocks and Implementation
In HiMSNet, the building blocks for temporal and spatial modeling are LSTM [35] and GCNConv [51] respectively.
LSTM [35] is a type of RNN with long-term modeling capabilities, and is widely used in sequential data analysis, such as time series. At each time step, LSTM [35] uses the current input $\mathbf{x}_{t}$, previous hidden state $\mathbf{h}_{t-1}$ and previous cell state $\mathbf{c}_{t-1}$ to compute the three internal gates $i$, $f$ and $g$, the output $\mathbf{o}_{t}$, new hidden state $\mathbf{h}_{t}$ and cell state $\mathbf{c}_{t}$. The three internal gates and the output have individual learnable parameters for input and hidden states. For example, in the $g$ gate $\mathbf{W}_{ig}$ is the weight matrix for the input, and $\mathbf{W}_{hg}$ is for hidden state (the notations for biases $\mathbf{b}$ are similar). The computation can be expressed using the set of formulas in Equation 7, where $\sigma$ is the sigmoid function, $\odot$ is the element-wise multiplication, and $\tanh$ is the hyperbolic tangent function.
$\mathbf{i}_{t}$ $=\sigma\left(\mathbf{W}_{ii}\mathbf{x}_{t}+\mathbf{b}_{ii}+\mathbf{W}_{hi}\mathbf{h}_{t-1}+\mathbf{b}_{hi}\right)$ (7)
$\mathbf{f}_{t}$ $=\sigma\left(\mathbf{W}_{if}\mathbf{x}_{t}+\mathbf{b}_{if}+\mathbf{W}_{hf}\mathbf{h}_{t-1}+\mathbf{b}_{hf}\right)$
$\mathbf{g}_{t}$ $=\tanh\left(\mathbf{W}_{ig}\mathbf{x}_{t}+\mathbf{b}_{ig}+\mathbf{W}_{hg}\mathbf{h}_{t-1}+\mathbf{b}_{hg}\right)$
$\mathbf{o}_{t}$ $=\sigma\left(\mathbf{W}_{io}\mathbf{x}_{t}+\mathbf{b}_{io}+\mathbf{W}_{ho}\mathbf{h}_{t-1}+\mathbf{b}_{ho}\right)$
$\mathbf{c}_{t}$ $=\mathbf{f}_{t}\odot\mathbf{c}_{t-1}+\mathbf{i}_{t}\odot\mathbf{g}_{t}$
$\mathbf{h}_{t}$ $=\mathbf{o}_{t}\odot\tanh\left(\mathbf{c}_{t}\right)$
GCNConv [51] is a graph neural network module for learning spatial correlation in the data. The feature for node $i$ at the $r$-th layer can be computed using the features of its neighbors $\mathcal{N}(i)$, including itself, by Equation 8:
$\mathbf{x}_{i}^{(r)}=\sum_{j\in\mathcal{N}(i)\cup\{i\}}\frac{1}{\sqrt{D(i)\cdot D(j)}}\cdot\left(\mathbf{W}\mathbf{x}_{j}^{(r-1)}\right),$ (8)
where $\mathbf{W}$ is the learnable parameters, $D(i)$ is the degree of node $i$ in the graph, and $\mathbf{x}_{j}^{r-1}$ is the feature of node $j$ at the previous layer.
To provide a reasonable baseline model, the implementation of HiMSNet is kept as concise as possible. The first step in HiMSNet is value embedding, i.e., the values and corresponding timestamps are embedded into a feature space using a linear layer, whose output dimension is 64. Specifically, the missing values, which appear when a road is not monitored or no vehicle passes a monitored road, are replaced with learnable neural network parameters of equal size. After the embedding, the drone modality has 2 layers of 1D convolution with kernel size 3 and stride 3 to down-sample the temporal resolution. Then, each modality has an individual temporal encoder $\mathcal{H}_{e}$ consisting of 3-layer LSTM [35] with hidden size 64, and their last step outputs will be used as the local feature $\mathbf{z}_{p}$. The decoders for both segment-level and regional predictions ($\mathcal{H}_{d}$ and $\mathcal{H}_{I}$) are 2-layer MLPs with doubled hidden size (128) and output dimension 10 to predict 10 future time steps (30 minutes). Contrary to modality-specific temporal encoders, the decoder is shared as the features have been combined. The message encoding layer $\mathcal{R}_{1}$ is a 1 $\times$ 1 convolution layer that projects the features while keeping the dimension unchanged. Accordingly, the message decoding layer $\mathcal{R}_{2}$ inherits the same structure. The GME module is implemented with 3 layers of GCNConv [51], each of them followed by ReLU activation and Layer Normalization [52].